Find the number of positive integral solution of abc=30 ?


i) 27 ii)81 iii)243


(Ans : 243)


plzz xplain !!!!!!

Grade:Upto college level

2 Answers

bhanuveer danduboyina
13 years ago

Use multinomial theorem

bhavin gupta
6 years ago
1*30*1=3 ways
2*5*3=3ways
6*5*1=3ways
15*2*1=3ways
10*3*1=3ways
 
so total=3*3*3*3*3=243
 
i hope this is right
this is the way i found the answer